# Non-commutative Iwasawa theory of abelian varieties over global function fields

### v1

@@ 4. Surjectivity of localisation map
$\mathrm{ker}\Big{(}H^{1}(G_{S}(L),T_{p}A)\to\prod_{w}H^{1}(L_{w},T_{p}A)\Big{)},$
where the product runs over all primes $w$ of $L$. We then define the compact Selmer group over $F_{\infty}$ by
@@
Define the continuous Iwasawa cohomology of $T_{p}A$ by
@@
where the inverse limit is taken over all finite Galois extensions $L$ of $F$ contained in $F_{\infty}$, via the corestriction maps. Note in particular that ${\mathfrak{S}}(A/F_{\infty})$ is a submodule of $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)$.
We now record the following result which will be a key ingredient for our eventual proof of the surjectivity.
@@
$H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)$ is a torsion $\Lambda(G)$-module. Moreover, if $\dim(G)\geq 2$ or if $A_{p^{\infty}}(F_{\infty})$ is finite, then $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)=0$.
@@
We consider the following spectral sequence of Jannsen (see [16, Theorem 1, Theorem 11]; see also [23, Theorem 4.5.1] and [27, Theorem 8.5.6]))
@@
The low degree terms corresponding to this spectral sequence give us the following exact sequence
(4.1) $0\to{\mathrm{Ext}}^{1}_{\Lambda(G)}\left(A_{p^{\infty}}(F_{\infty})^{\vee},\Lambda(G)\right)\to H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)\to{\mathrm{Ext}}^{0}_{\Lambda(G)}\left(H^{1}\left(G_{S}(F_{\infty}),A_{p^{\infty}})\right)^{\vee},\Lambda(G)\right).$
By Lemma 3.3, we have ${\mathrm{rank}}_{\Lambda(G)}\left(H^{1}\left(G_{S}(F_{\infty}),A_{p^{\infty}})\right)^{\vee}\right)=0$, so the right-most term vanishes. Since ${\mathrm{Ext}}^{1}_{\Lambda(G)}\left(A_{p^{\infty}}(F_{\infty})^{\vee},\Lambda(G)\right)$ is a torsion $\Lambda(G)$-module (cf.  [37, Proposition 3.5 (iii) (a)]), it follows that $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)$ is a torsion $\Lambda(G)$-module. Now, $A_{p^{\infty}}(F_{\infty})^{\vee}$ is finitely generated over ${\mathbb{Z}}_{p}$, so the assumption $\dim(G)\geq 2$ tells us that $A_{p^{\infty}}(F_{\infty})^{\vee}$ is a pseudo-null module over $\Lambda(G)$, and the left-most term of (4.2) also vanishes. Thus, $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)=0$, as required. ∎
@@
Let $A$ be an abelian variety defined over a global function field $F$ of characteristic prime to $p$. Let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Then the localisation map $\lambda_{S}(A/F_{\infty})$ is surjective, giving rise to the following short exact sequence:
$0\to{\mathrm{Sel}}(A/F_{\infty})\to H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\stackrel{{\scriptstyle\lambda_{S}(A/F_{\infty})}}{{\longrightarrow}}\bigoplus_{v\in S}J_{v}(A/F_{\infty})\to 0$
@@
As the conclusion of the theorem remains unchanged upon replacing the base field field $F$ by a larger field contained in $F_{\infty}$, we may assume that $G$ is pro-$p$. The Poitou-Tate sequence ([30, Proposition A.3.2]) gives an exact sequence
@@
where the final ``$0"$ term is a consequence of Lemma 3.2. It therefore remains to show that ${\mathfrak{S}}(A^{*}/F_{\infty})=0$.
We consider the following spectral sequence of Jannsen (see [16, Theorem 1, Theorem 11]; see also [23, Theorem 4.5.1] and [27, Theorem 8.5.6]))
$E_{2}^{i,j}={\mathrm{Ext}}^{i}_{\Lambda(G)}\left(H^{j}\left(G_{S}(F_{\infty}),A^{*}_{p^{\infty}})\right)^{\vee},\Lambda(G)\right)\Rightarrow H^{i+j}_{\mathrm{Iw}}(F_{\infty},T_{p}A^{*}).$
The low degree terms corresponding to this spectral sequence give us the following exact sequence
(4.2) $0\to{\mathrm{Ext}}^{1}_{\Lambda(G)}\left(A^{*}_{p^{\infty}}(F_{\infty})^{\vee},\Lambda(G)\right)\to H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A^{*})\to{\mathrm{Ext}}^{0}_{\Lambda(G)}\left(H^{1}\left(G_{S}(F_{\infty}),A^{*}_{p^{\infty}})\right)^{\vee},\Lambda(G)\right).$
By Lemma 3.3, we have ${\mathrm{rank}}_{\Lambda(G)}\left(H^{1}\left(G_{S}(F_{\infty}),A_{p^{\infty}})\right)^{\vee}\right)=0$, so the right-most term vanishes. Now, $A_{p^{\infty}}(F_{\infty})^{\vee}$ is finitely generated over ${\mathbb{Z}}_{p}$. Therefore, if $\dim(G)\geq 2$, then $A_{p^{\infty}}(F_{\infty})^{\vee}$ is a pseudo-null module over $\Lambda(G)$, and so the left-most term of (4.2) also vanishes. Thus, we have shown that ${\mathfrak{S}}(A^{*}/F_{\infty})=0$ when $\dim(G)\geq 2$.
It therefore remains to consider the case where $F_{\infty}=F^{\mathrm{cyc}}$. This case is slightly more involved, owing to the fact that we do not know the finiteness of $A_{p^{\infty}}(F^{\mathrm{cyc}})$ in general (but see Remark 4.3 in the case where $A$ is an elliptic curve without complex multiplication). By base changing $F$, if necessary, we may assume that every prime of $S$ is inert in $F^{\mathrm{cyc}}/F$. Thus, by abuse of notation, given $v\in S$, we also denote by $v$ the unique prime of $F^{\mathrm{cyc}}$ lying above $v$, and identify the Galois groups ${\mathrm{Gal}}(F^{\mathrm{cyc}}/F)$ with ${\mathrm{Gal}}(F^{\mathrm{cyc}}_{v}/F_{v})$, denoting them both by $\Gamma$. For each $v\in S$, we consider the following local version of Jannsen’s spectral sequence (see [16, Theorem 11], [23, Theorem 4.2.2] or [27, Theorem 5.2.6]) for the dual abelian variety $A^{*}$:
$E_{2}^{i,j}={\mathrm{Ext}}^{i}_{\Lambda(\Gamma)}\left(H^{j}\left(F^{\mathrm{cyc}}_{v},A_{p^{\infty}}^{*})\right)^{\vee},\Lambda(\Gamma)\right)\Rightarrow H^{i+j}_{\mathrm{Iw}}(F^{\mathrm{cyc}}_{v},T_{p}A^{*}).$
To simplify notation, we write $\mathrm{E}^{i}(-)={\mathrm{Ext}}^{i}_{\Lambda(\Gamma)}\left(-,\Lambda(\Gamma)\right)$. Analogously to the global case, the low degree terms of the above spectral sequence give an exact sequence
@@
The global and local spectral sequences satisfy a functoriality property [23, Theorem 4.5.1], yielding the following commutative diagram:
[…]
where the vertical maps are induced by the natural localisation maps. Note that $\ker g={\mathfrak{S}}(A^{*}/F^{\mathrm{cyc}})$. Thus, it remains to show that $f$ is injective. Since $S$ is assumed to be non-empty, writing $f=\oplus_{v\in S}f_{v}$, it suffices to show that
@@
is injective for each $v\in S$. But since both $A^{*}_{p^{\infty}}(F^{\mathrm{cyc}})^{\vee}$ and $A^{*}_{p^{\infty}}(F^{\mathrm{cyc}}_{v})^{\vee}$ are finitely generated over ${\mathbb{Z}}_{p}$, it follows by [28, Corollary 5.5.7] that $f_{v}$ is given by
@@
which is injective since it is induced by the inclusion $A^{*}_{p^{\infty}}(F^{\mathrm{cyc}})\hookrightarrow A^{*}_{p^{\infty}}(F^{\mathrm{cyc}}_{v})$. This completes the proof of the theorem. ∎
@@
In the case where $A=E$ is an elliptic curve without complex multiplication, the proof of Theorem 4.2 can be simplified. By a theorem of Deuring, this is equivalent to saying that $E$ is not defined over a finite field; equivalently, $E$ is non-isotrivial. In particular, we do not have to separately consider the case $F_{\infty}=F^{\mathrm{cyc}}$, since Lemma 4.1 holds without the assumption $\dim(G)\geq 2$. One way to see this is the result of Ulmer, which says that $E(\bar{k}F)$ is a finitely generated abelian group (see [35, §6.3]). Since $F^{\rm cyc}$ is a subfield of $\bar{k}F$, it follows that $E(F^{\rm cyc})$ is a finitely generated abelian group, and thus $E_{p^{\infty}}(F^{\rm cyc})$ is finite. It follows that $E_{p^{\infty}}(F^{\rm cyc})$ is pseudo-null over $\Lambda(G)$, and we always have $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}E)$ injecting into the ${\mathrm{Ext}}^{0}$-term.
@@ 5. Akashi series
Denote by $\mathcal{Q}(\Gamma)$ the field of fractions of $\Lambda(\Gamma)$. To each $M$ in ${\mathfrak{M}}_{H}(G)$, we now attach a non-zero element $\mathrm{Ak}(M)$ of $\mathcal{Q}(\Gamma)$, which is uniquely determined up to multiplication by a unit in $\Lambda(\Gamma)$. It can be shown (see [8, Lemma 3.1]) that the homology groups
@@
are finitely generated torsion $\Lambda(\Gamma)$-modules. We fix from now on a topological generator of $\Gamma$, and identify $\Lambda(\Gamma)$ with the formal power series ring ${\mathbb{Z}}_{p}[[T]]$ by mapping this topological generator to $1+T$. Let $f_{M,i}\in\Lambda(\Gamma)$ be a characteristic power series for $H_{i}(H,M)$, and define
@@
This product is finite because $H_{i}(H,M)=0$ for $i\geq\dim(H)$, and it is well defined up to multiplication by a unit in $\Lambda(\Gamma)$ because each $f_{M,i}$ is. We call $\mathrm{Ak}(M)$ the Akashi series of $M$. If $Y$ is a discrete $p$-primary $G$-module such that $M=Y^{\vee}\in{\mathfrak{M}}_{H}(G)$, then we call $\mathrm{Ak}(M)$ also the Akashi series of $Y$, and by abuse of notation, we also denote it by $\mathrm{Ak}(Y)$. Via the duality
@@
we will work with $H$-cohomology groups in our computation of Akashi series without further mention. It is indeed a generalisation of the notion of characteristic power series, because if $M$ is a finitely generated torsion $\Lambda(\Gamma)$-module (we can view $H=\{1\}$), then the Akashi series of $M$ is exactly the $\Lambda(\Gamma)$-characteristic power series of $M$. In this case, we simply denote $\mathrm{Ak}(M)$ by $f_{M}$, or $f_{Y}$, again by abuse of notation.
@@
Given an exact sequence
@@
of modules in the category ${\mathfrak{M}}_{H}(G)$, their Akashi series satisfy $\mathrm{Ak}(M_{2})=\mathrm{Ak}(M_{1})\cdot\mathrm{Ak}(M_{3})$.
@@
This follows from considering the long exact sequence of $H$-homology derived from (5.2) and applying the multiplicativity of $\Lambda(\Gamma)$-characteristic series in short exact sequences. ∎
@@
We note that the original definition of ${\mathfrak{M}}_{H}(G)$ in [9] differs from the definition in the subsequent paper [8] or in this paper. One consequence of this is that it is no longer true that if $M$ is a module in ${\mathfrak{M}}_{H}(G)$ then $\mathrm{Ak}(M)=\mathrm{Ak}(M/M(p))$, as shown in [9, Lemma 4.1]. This is because $M(p)$ need not be finitely generated over $\Lambda(H)$, and so its homology group may not be finitely generated over ${\mathbb{Z}}_{p}$. Therefore, even though its homology group is annihilated by some power of $p$, we cannot conclude that they are finite. In general, $\mathrm{Ak}(M(p))$ can be non-trivial. In case $G$ is pro-$p$ without $p$-torsion, it is given by $\mathrm{Ak}(M(p))=p^{\mu(M)}$ (see [1, 1.6]), where $\mu(M)$ denotes the $\mu$-invariant of $M$ defined in [37, 4].
@@
$0\to{\mathrm{Sel}}(A/F_{\infty})\to H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\to\bigoplus_{v\in S}J_{v}(A/F_{\infty})\to 0$
Since ${\mathrm{Sel}}(A/F_{\infty})$ and $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})$ are cofinitely generated $\Lambda(H)$-modules by Lemma 3.3 and Theorem 3.4, so is $\bigoplus_{v\in S}J_{v}(A/F_{\infty})$. Therefore, we may apply Lemma 5.1 to conclude that
@@
Therefore, the determination of $\mathrm{Ak}\Big{(}{\mathrm{Sel}}(A/F_{\infty})\Big{)}$ is reduced to computing $\mathrm{Ak}\Big{(}H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\Big{)}$ and $\mathrm{Ak}\left(\bigoplus_{v\in S}J_{v}(A/F_{\infty})\right)$. These will be the content of the next two lemmas.
@@
$\mathrm{Ak}\Big{(}H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\Big{)}=f_{H^{1}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})}\prod_{i\geq 1}\left(f_{H^{i}\left(H,A_{p^{\infty}}(F_{\infty})\right)}\right)^{(-1)^{i}}$.
@@
degenerates to yield an exact sequence
@@
for $i\geq 1$. The conclusion now follows from a direct calculation using (5.4) and (5.5). ∎
Since $\mathrm{char}(F_{v})\neq p$ for each $v\in S$, it follows from a theorem of Iwasawa [28, Theorem 7.5.3] that the decomposition subgroup $G_{w}$ of $w$ in $G$ has dimension at most $2$. Let $S^{\prime}$ denote the subset of $S$ consisting of primes whose inertia group in $G$ is infinite. We note that this is equivalent to requiring that the decomposition subgroup $G_{w}$ has dimension exactly $2$. Indeed, since the residue characteristic is prime to $p$ and $G$ is a $p$-adic Lie group, the wild inertia subgroup is finite. Moreover, the maximal unramified $p$-extension of $F_{v}$ is $F_{w}^{\mathrm{cyc}}$, and the maximal tamely ramified extension of $F_{v}$ is a topologically cyclic extension over the maximal unramified extension.
@@
We have $\mathrm{Ak}\left(\bigoplus_{v\in S}J_{v}(A/F_{\infty})\right)=\prod_{v\in S\setminus S^{\prime}}f_{J_{v}(A/F^{\mathrm{cyc}})}$

### v2

@@ 4. Surjectivity of localisation map
$\mathrm{ker}\Big(H^{1}(G_{S}(L),T_{p}A)\to\bigoplus_{w\in S}H^{1}(L_{w},T_{p}A)\Big),$
where the sum runs over all primes $w$ of $L$ lying above the set $S$. We then define the compact Selmer group over $F_{\infty}$ as the inverse limit
@@
where the limit is taken over all finite extensions $L$ of $F$ contained in $F_{\infty}$. We also define the continuous Iwasawa cohomology of $T_{p}A$ by
@@
where the limit is taken over all finite Galois extensions with transition maps given by the corestriction maps on cohomology. In particular, ${\mathfrak{S}}(A/F_{\infty})$ is a submodule of $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)$.
We now record the following result, which will be a key ingredient in our eventual proof of surjectivity.
@@
The module $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)$ is a torsion $\Lambda(G)$-module. Moreover, if $\dim(G)\geq 2$ or if $A_{p^{\infty}}(F_{\infty})$ is finite, then $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)=0$.
@@
We consider the following spectral sequence of Jannsen (see [20, Theorem 1, Theorem 11]; see also [30, Theorem 4.5.1] and [34, Theorem 8.5.6])
@@
The low-degree terms of this spectral sequence yield the following exact sequence
(4.1) $0\to{\mathrm{Ext}}^{1}_{\Lambda(G)}\left(A_{p^{\infty}}(F_{\infty})^{\vee},\Lambda(G)\right)\to H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)\to{\rm Hom}_{\Lambda(G)}\left(H^{1}\left(G_{S}(F_{\infty}),A_{p^{\infty}})\right)^{\vee},\Lambda(G)\right).$
By Lemma 3.3, we have ${\mathrm{rank}}_{\Lambda(G)}\left(H^{1}\left(G_{S}(F_{\infty}),A_{p^{\infty}})\right)^{\vee}\right)=0$, so the right-most term vanishes. Since ${\mathrm{Ext}}^{1}_{\Lambda(G)}\left(A_{p^{\infty}}(F_{\infty})^{\vee},\Lambda(G)\right)$ is a torsion $\Lambda(G)$-module (cf.  [46, Proposition 3.5 (iii) (a)]), it follows that $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)$ is a torsion $\Lambda(G)$-module. Now by an observation of Jannsen (cf. [19, Corollary 2.6(b)(c)]), if $\dim(G)\geq 2$ or if $A_{p^{\infty}}(F_{\infty})$ is finite, then ${\mathrm{Ext}}^{1}_{\Lambda(G)}\left(A_{p^{\infty}}(F_{\infty})^{\vee},\Lambda(G)\right)=0$. Therefore, both the left-most and right-most terms in (4.1) vanish, and we conclude that $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A)=0$, as required. ∎
@@
Let $A$ be an abelian variety defined over a global function field $F$ of characteristic prime to $p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Then the localisation map $\lambda_{S}(A/F_{\infty})$ is surjective, yielding the following short exact sequence:
(4.2) $0\longrightarrow\operatorname{Sel}(A/F_{\infty})\longrightarrow H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\overset{\lambda_{S}(A/F_{\infty})}{\xrightarrow{\hskip 85.35826pt}}\bigoplus_{v\in S}J_{v}(A/F_{\infty})\longrightarrow 0.$
@@
Since the conclusion remains unchanged upon replacing the base field $F$ by a larger field contained in $F_{\infty}$, we may assume that $G$ is pro-$p$. The Poitou–Tate sequence (see [35, Proof of Theorem 8.7.9]) gives the exact sequence
@@
where the final “0” term follows from Lemma 3.2. Thus, it suffices to show that ${\mathfrak{S}}(A^{*}/F_{\infty})=0$. By definition, ${\mathfrak{S}}(A^{*}/F_{\infty})$ is contained in $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A^{*})$. If ${\rm dim}(G)\geq 2$, then Lemma 4.1 implies that $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A^{*})=0$, and hence ${\mathfrak{S}}(A^{*}/F_{\infty})=0$.
It remains to consider the case $F_{\infty}=F^{\mathrm{cyc}}$. This case is slightly more delicate, as we do not know a priori whether $A_{p^{\infty}}(F^{\mathrm{cyc}})$ is finite (see Remark 4.3 for the elliptic curve case). By replacing $F$ with a finite extension if necessary, we may assume that every prime in $S$ remains inert in $F^{\mathrm{cyc}}/F$. For each $v\in S$, we denote by $v$ also the unique prime of $F^{\mathrm{cyc}}$ lying above $v$, and identify the Galois groups $\Gamma={\mathrm{Gal}}(F^{\mathrm{cyc}}/F)$ with ${\mathrm{Gal}}(F^{\mathrm{cyc}}_{v}/F_{v})$. On the global side, the Jannsen spectral sequence (see [20, Theorem 1, Theorem 11]; see also [30, Theorem 4.5.1] and [34, Theorem 8.5.6])) gives
${\mathrm{Ext}}^{i}_{\Lambda(\Gamma)}\left(H^{j}\left(G_{S}(F^{\mathrm{cyc}}),A^{*}_{p^{\infty}})\right)^{\vee},\Lambda(\Gamma)\right)\Rightarrow H^{i+j}_{\mathrm{Iw}}(F^{\mathrm{cyc}},T_{p}A^{*}).$
The low-degree terms yield the following exact sequence
(4.3) $0\to{\mathrm{Ext}}^{1}_{\Lambda(\Gamma)}\left(A^{*}_{p^{\infty}}(F^{\mathrm{cyc}})^{\vee},\Lambda(G)\right)\to H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}A^{*})\to{\mathrm{Ext}}^{0}_{\Lambda(\Gamma)}\left(H^{1}\left(G_{S}(F^{\mathrm{cyc}}),A^{*}_{p^{\infty}})\right)^{\vee},\Lambda(\Gamma)\right).$
By Lemma 3.3, we have ${\mathrm{rank}}_{\Lambda(\Gamma)}\left(H^{1}\left(G_{S}(F_{\infty}),A_{p^{\infty}})\right)^{\vee}\right)=0$, so the right-most term vanishes.
For each $v\in S$, there is also a local version of Jannsen’s spectral sequence (see [20, Theorem 11], [30, Theorem 4.2.2] or [34, Theorem 5.2.6]) for $A^{*}$:
${\mathrm{Ext}}^{i}_{\Lambda(\Gamma)}\left(H^{j}\left(F^{\mathrm{cyc}}_{v},A_{p^{\infty}}^{*})\right)^{\vee},\Lambda(\Gamma)\right)\Rightarrow H^{i+j}_{\mathrm{Iw}}(F^{\mathrm{cyc}}_{v},T_{p}A^{*}).$
Let us simplify notation by writing $\mathrm{E}^{i}(-)={\mathrm{Ext}}^{i}_{\Lambda(\Gamma)}\left(-,\Lambda(\Gamma)\right)$. The low-degree terms give an exact sequence:
@@
The global and local spectral sequences are functorial [30, Theorem 4.5.1], which yields the following commutative diagram with exact rows:
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{E}^{1}\left(A^{*}_{p^{\infty}}(F^{\mathrm{cyc}})^{\vee}\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${H^{1}_{\mathrm{Iw}}(F^{\mathrm{cyc}},T_{p}A^{*})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{v\in S}\mathrm{E}^{1}\left(A^{*}_{p^{\infty}}(F^{\mathrm{cyc}}_{v})^{\vee}\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{v\in S}H^{1}_{\mathrm{Iw}}(F^{\mathrm{cyc}}_{v},T_{p}A^{*})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{v\in S}\mathrm{E}^{0}\left(H^{1}(F^{\mathrm{cyc}}_{v},A^{*}_{p^{\infty}})^{\vee}\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0,}$
where the vertical maps are induced by the natural localisation maps. Note that $\ker g={\mathfrak{S}}(A^{*}/F^{\mathrm{cyc}})$. So it remains to show that $f$ is injective. Since $S$ is non-empty, and $f=\bigoplus_{v\in S}f_{v}$, it suffices to show that each
@@
is injective. But since both $A^{*}_{p^{\infty}}(F^{\mathrm{cyc}})^{\vee}$ and $A^{*}_{p^{\infty}}(F^{\mathrm{cyc}}_{v})^{\vee}$ are finitely generated ${\mathbb{Z}}_{p}$-modules, [35, Corollary 5.5.7] implies that $f_{v}$ is given by
@@
which is injective, as it is induced by the inclusion $A^{*}_{p^{\infty}}(F^{\mathrm{cyc}})\hookrightarrow A^{*}_{p^{\infty}}(F^{\mathrm{cyc}}_{v})$. This completes the proof. ∎
@@
In the case where $A=E$ is an elliptic curve without complex multiplication, the proof of Theorem 4.2 can be simplified. By a theorem of Deuring, this is equivalent to the condition that $E$ is not defined over a finite field; equivalently, $E$ is non-isotrivial. In particular, we do not need to treat the case $F_{\infty}=F^{\mathrm{cyc}}$ separately, since Lemma 4.1 holds without requiring the assumption $\dim(G)\geq 2$. One way to see this is via a result of Ulmer, which asserts that $E(\bar{k}F)$ is a finitely generated abelian group (see [44, §6.3]). Since $F^{\rm cyc}$ is a subfield of $\bar{k}F$, it follows that $E(F^{\rm cyc})$ is a finitely generated abelian group. Consequently, $E_{p^{\infty}}(F^{\rm cyc})$ is finite. It implies that $E_{p^{\infty}}(F^{\rm cyc})$ is pseudo-null as a $\Lambda(G)$-module, and therefore $H^{1}_{\mathrm{Iw}}(F_{\infty},T_{p}E)$ injects into the ${\mathrm{Ext}}^{0}$-term of the corresponding spectral sequence.
@@ 5. Akashi series
Let $\mathcal{Q}(\Gamma)$ denote the field of fractions of $\Lambda(\Gamma)$. To each $M\in{\mathfrak{M}}_{H}(G)$, we associate a non-zero element $\mathrm{Ak}(M)\in\mathcal{Q}(\Gamma)$, which is uniquely determined up to multiplication by a unit in $\Lambda(\Gamma)$. It can be shown (see [9, Lemma 3.1]) that the homology groups
@@
are finitely generated torsion $\Lambda(\Gamma)$-modules. We fix, from now on, a topological generator of $\Gamma$, and identify $\Lambda(\Gamma)$ with the formal power series ring ${\mathbb{Z}}_{p}[[T]]$ by mapping this topological generator to $1+T$. Let $f_{M,i}\in\Lambda(\Gamma)$ denote a characteristic power series of $H_{i}(H,M)$, and define
@@
This product is finite because $H_{i}(H,M)=0$ for $i\geq\dim(H)$, and it is well defined up to multiplication by a unit in $\Lambda(\Gamma)$, since each $f_{M,i}$ is. We call $\mathrm{Ak}(M)$ the Akashi series of $M$. If $Y$ is a discrete $p$-primary $G$-module such that $M=Y^{\vee}\in{\mathfrak{M}}_{H}(G)$, then we also refer to $\mathrm{Ak}(M)$ as the Akashi series of $Y$, and by abuse of notation, write it as $\mathrm{Ak}(Y)$. Via the duality
@@
we will henceforth work with $H$-cohomology groups in computating the Akashi series, without further mention. This definition generalises the notion of the characteristic power series. Specifically, if $M$ is a finitely generated torsion $\Lambda(\Gamma)$-module (i.e. when $H=\{1\}$), then the Akashi series of $M$ coincides with the $\Lambda(\Gamma)$-characteristic power series of $M$, and we denote $\mathrm{Ak}(M)$ simply by $f_{M}$, or $f_{Y}$.
@@
Given a short exact sequence
@@
in the category ${\mathfrak{M}}_{H}(G)$, the Akashi series satisfy $\mathrm{Ak}(M_{2})=\mathrm{Ak}(M_{1})\cdot\mathrm{Ak}(M_{3})$.
@@
This follows by considering the long exact sequence in $H$-homology derived from (5.2), and applying the multiplicativity of $\Lambda(\Gamma)$-characteristic series in a short exact sequence. ∎
@@
We note that the original definition of ${\mathfrak{M}}_{H}(G)$ in [10] differs from the one used in [9] and in this paper. One consequence of this discrepancy is that, in general, $\mathrm{Ak}(M)\neq\mathrm{Ak}(M/M(p))$, as shown in [10, Lemma 4.1]. This is because $M(p)$ need not be finitely generated over $\Lambda(H)$, and hence its homology group may not be finitely generated over ${\mathbb{Z}}_{p}$. Therefore, even though its homology group is annihilated by some power of $p$, we cannot conclude that they are finite. In general, $\mathrm{Ak}(M(p))$ may be non-trivial. In the case where $G$ is a pro-$p$ group without $p$-torsion, we have $\mathrm{Ak}(M(p))=p^{\mu(M)}$ (see [1, 1.6]), where $\mu(M)$ denotes the generalised $\mu$-invariant of $M$, as defined in [18, 46, 4].
@@
$0\to{\mathrm{Sel}}(A/F_{\infty})\to H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\to\bigoplus_{v\in S}J_{v}(A/F_{\infty})\to 0.$
Since ${\mathrm{Sel}}(A/F_{\infty})$ and $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})$ are cofinitely generated $\Lambda(H)$-modules by Lemma 3.3 and Theorem 3.4, so is $\bigoplus_{v\in S}J_{v}(A/F_{\infty})$. Hence, we may apply Lemma 5.1 to deduce
@@
Thus, determining $\mathrm{Ak}\Big({\mathrm{Sel}}(A/F_{\infty})\Big)$ reduces to computing
$\mathrm{Ak}\Big(H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\Big)\quad\textrm{and}\quad\mathrm{Ak}\left(\bigoplus_{v\in S}J_{v}(A/F_{\infty})\right),$
which are the subjects of the next two lemmas.
@@
$\mathrm{Ak}\Big(H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})\Big)=f_{H^{1}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})}\cdot\prod_{i\geq 1}\left(f_{H^{i}\left(H,A_{p^{\infty}}(F_{\infty})\right)}\right)^{(-1)^{i}}$.
@@
degenerates and yields the following exact sequence
@@
for $i\geq 1$. The lemma now follows from a direct calculation using (5.4) and (5.5). ∎
Since $\mathrm{char}(F_{v})\neq p$ for each $v\in S$, it follows from a theorem of Iwasawa [35, Theorem 7.5.3] that the decomposition subgroup $G_{w}$ of any prime $w$ of $F_{\infty}$ lying above $v$ has dimension at most $2$. Let $S^{\prime}$ denote the subset of $S$ consisting of primes whose inertia group in $G$ is infinite. This is equivalent to requiring that the decomposition subgroup $G_{w}$ has dimension exactly $2$. Indeed, since the residue characteristic is prime to $p$ and $G$ is a $p$-adic Lie group, the wild inertia subgroup is finite. Moreover, the maximal unramified $p$-extension of $F_{v}$ is $F_{v}^{\mathrm{cyc}}$, and the maximal tamely ramified extension of $F_{v}$ is a topologically cyclic extension over the maximal unramified extension.
@@
$\mathrm{Ak}\left(\bigoplus_{v\in S}J_{v}(A/F_{\infty})\right)=\prod_{v\in S\setminus S^{\prime}}f_{J_{v}(A/F^{\mathrm{cyc}})}$